# Spectrum-Aware Debiasing: A Modern Inference Framework with Applications to Principal Components Regression

### v4

@@ 3.1 Well-definedness of our debiasing formula
Here $h^{\prime\prime}(\cdot)$ refers to the extended version we defined using Lemma 2.7.
The following Proposition is restated from Proposition B.10 in Appendix.
@@
admits a unique solution in $(0,+\infty)$ if and only if there exists some $i\in[p]$ such that $h^{\prime\prime}(\hat{{\beta}}_{i})\neq+\infty$ and at least one of the following holds: (i) $\left\|h^{\prime\prime}(\hat{\bm{\beta}})\right\|_{0}=p$; (ii) $\mathbf{X}^{\top}\mathbf{X}$ is non-singular; (iii) $\norm{d}_{0}+\norm{h^{\prime\prime}(\hat{\bm{\beta}})}_{0}>p$
Example 3.2.
As a concrete example, the assumptions of Proposition 3.1 hold for the Elastic Net with $\lambda_{2}>0$.
@@ 3.2 The procedure
$\hat{\bm{\beta}}^{u}=\hat{\bm{\beta}}+\widehat{\mathsf{adj}}^{-1}\mathbf{X}^{\top}(\mathbf{y}-\mathbf{X}\hat{\bm{\beta}}).$ (25)
@@ 3.3 Asymptotic normality
Theorem 3.1 below states that the empirical distribution of $(\hat{\tau}_{*}^{-1/2}(\hat{{\beta}}^{u}_{i}-{\beta}^{\star}_{i}))_{i=1}^{p}$ converges to a standard Gaussian.
@@
Next, we describe a consistent estimator for the asymptotic variance $\tau_{*}$. We require some intermediate quantities that depend on the observed data and the choice of the penalty. We define these next. Later in Section 5, we will provide intuition as to why these intermediate quantities are important and how we construct the variance estimator.
@@
$\hat{\eta}_{*}(\mathbf{X},\mathbf{y},h)\leftarrow\left(\frac{1}{p}{{\sum}}_{j=1}^{p}\frac{1}{\widehat{\mathsf{adj}}+h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)}\right)^{-1}$ (26)
$\hat{\tau}_{**}(\mathbf{X},\mathbf{y},h)\leftarrow\frac{\left\|\quantity(\mathbf{I}_{n}+\frac{1}{\hat{\eta}_{*}-\widehat{\mathsf{adj}}}\mathbf{X}\mathbf{X}^{\top})\quantity(\mathbf{y}-\mathbf{X}\hat{\bm{\beta}})\right\|^{2}-n\hat{\sigma}^{2}}{\sum_{i=1}^{p}d_{i}^{2}}$
$\hat{\tau}_{*}(\mathbf{X},\mathbf{y},h)\leftarrow\frac{1}{p}{{\sum}}_{i=1}^{p}\frac{\hat{\eta}_{*}^{2}d_{i}^{2}\hat{\sigma}^{2}+\quantity(d_{i}^{2}-\widehat{\mathsf{adj}}+2\hat{\eta}_{*})\quantity(\widehat{\mathsf{adj}}-d_{i}^{2})\quantity(\hat{\eta}_{*}-\widehat{\mathsf{adj}})^{2}\hat{\tau}_{**}}{\quantity(d_{i}^{2}-\widehat{\mathsf{adj}}+\hat{\eta}_{*})^{2}\quantity(\widehat{\mathsf{adj}})^{2}}$
@@
We illustrate Theorem 3.1 in Figure 1 under five different right-rotationally-invariant designs (cf. Remark E.2 in Appendix) with non-trivial correlation structures, and compare with Degrees-of-Freedom Debiasing with $\mathbf{M}=\mathbf{I}_{p}$. The corresponding QQ-plot can be found in Figure 5 in Appendix. We observe that our method outperforms Degrees-of-Freedom Debiasing by a margin.
@@
We call a sequence of random variables $\left(\mathsf{V}_{i}\right)_{i=1}^{p}$ exchangeable if for any permutation $\pi$ of the indices $1,...,p$, the joint distribution of the permuted sequence $\left(\mathsf{V}_{\pi(i)}\right)_{i=1}^{p}$ is the same as the original sequence.
Corollary 3.8 below is a consequence of Theorem 3.1. We defer its proof to Section B.5 in Appendix. A numerical demonstration is included in Section E.3 from Appendix.
@@ 3.4 Inference
Figure 2: The above plots the TPR and FPR of the hypothesis testing procedure defined in (28) with significance level $\alpha$ and FCP of the constructed confidence intervals (29) with $b=\Phi^{-1}(1-\alpha/2),a=\Phi^{-1}(\alpha/2)$ as $\alpha$ on the x-axis varies from 00 to $1$, for both degrees-of-freedom ($\mathsf{DF}$, blue) adjustment and Spectrum-Aware ($\mathsf{SA}$, red) adjustment. The setting here is the same as in Figure 1.
@@
for any $p\geq 1$. Theorem 3.1 directly yield guarantees on the FPR, TPR and FCP as shown in Corollary 3.9 below. We defer the proof to Section B.6 in Appendix.
@@
    (|$\mathbb{P}\left(\absolutevalue{\mathsf{B}^{\star}}\in\left(\mu_{0},+\infty\right)\cup\{0\}\right)=1.$
    Then for any fixed $i$ such that ${\beta}^{\star}_{i}=0$, we have $\lim_{p\rightarrow\infty}\mathbb{P}\left(T_{i}=1\right)=\alpha,$ and the false positive rate satisfies that almost surely $\lim_{p\to\infty}\mathsf{FPR}(p)=\alpha.$ Refer also to Remark B.14 from Appendix for the exact asymptotic limit of TPR.
  • (b
@@ 4.1 Outliers and PC alignment
The assumptions made in our previous section exclude important scenarios where the design may contain outlier eigenvalues or the signal may align with an eigenvector of the sample covariance matrix. To address this challenge, we propose an enhanced Spectrum-Aware procedure whic relaxes Assumptions 2 and 3 to Assumption 8 below. To this end, denote $\mathcal{N}:=\left\{i\in[p]:d_{i}^{2}>0\right\},N:=|\mathcal{N}|$. We let $\mathcal{J}$ be a user-chosen, finite index set $\mathcal{J}\subseteq\mathcal{N}$ that contains $\mathcal{J}_{1}\cup\mathcal{J}_{2}$ as subset (See Remark 4.2). We denote its size as $J:=\absolutevalue{\mathcal{J}}$ |.
@@
where we used $\mathcal{J}(i)$ to denote the $i$-th index in $\mathcal{J}$. Both $\bm{\upsilon}^{\star}$ and $\bm{\zeta}^{\star}$ are unknown, and they can be either deterministic or random independent of $\mathbf{O},\bm{\varepsilon}$. If $\bm{\zeta}^{\star}$ is deterministic, we assume that $\bm{\zeta}^{\star}\stackrel{{\scriptstyle W_{2}}}{{\to}}\mathsf{C}^{\star}$ as $n,p\to\infty$, where $\mathsf{C}^{\star}$ is a random variable with finite variance. If $\bm{\zeta}^{\star}$ is random, we assume the same convergence holds almost surely. Furthermore, we assume that Assumption 2 holds except that, instead of (13), we only require $\lim_{p\to\infty}\max_{i\in[p]\setminus\mathcal{J}}d_{i}^{2}<+\infty.$
Under Assumption 8, $\bm{\beta}^{\star}_{\mathsf{al}}$ is the alignment component that aligns to $\mathcal{J}(i)$-th Principal Component (PC) $\mathbf{o}_{\mathcal{J}(i)}$ if the corresponding $\upsilon^{\star}_{i}$ is non-zero, while $\bm{\zeta}^{\star}$ is the complement component that is independent of the design. Note that when $\mathcal{J}=\emptyset$, Assumption 8 reduces to Assumptions 2 and 3 precisely.
@@
$\mathcal{J}$ needs to be a finite index set that contains both $\mathcal{J}_{1}$ and $\mathcal{J}_{2}$ as subsets. The index set $\mathcal{J}_{2}$ can be determined by observing the spectrum of $\mathbf{X}^{\top}\mathbf{X}$. $\mathcal{J}_{1}$ is generally not observed and requires prior information. However, we remark that eigenvectors indexed by $\mathcal{J}_{1}\cap\mathcal{J}_{2}$ tend to distort the debiasing procedure most severely. So often just including $\mathcal{J}_{2}$ (i.e. top few PCs) in $\mathcal{J}$ can significantly improve inference.
@@ 4.2 The PCR algorithms
Figure 3: top-left: Same setting as Figure 1 except for specific changes to the design distribution parameters that lead to more challenging scenarios (see details of $\mathsf{MatrixNormal}$-$\mathsf{B}$,…,$\mathsf{MultiCauchy}$ in Definition E.4 from Appendix). Rows 1–4 correspond to: (i) DF: Degrees-of-Freedom Debiasing as described in [12]; (ii) SA: Spectrum-Aware Debiasing as described in Definition 3.4; (iii): PCRDF: PCR-Degrees-of-Freedom Debiasing, that is, the procedure obtained from Section 4.3.3 on substituting Degrees-of-Freedom Debiasing for Spectrum-Aware Debiasing in the complement PCR step; (iv) PCRSA: PCR-Spectrum-Aware Debiasing as described in Section 4.3.3. The true signals $\bm{\beta}^{\star}$ for these designs are generated such that they align with the 2nd, 4th, and 6th PCs. Concretely, we generate the signal as follows: $\bm{\beta}^{\star}=\bm{\beta}^{\star}_{\mathsf{al}}+\bm{\zeta}^{\star}$ where the components of ${\zeta}^{\star}_{i}$ are i.i.d. draws satisfying ${\zeta}^{\star}_{i}\sim 0.24\cdot N(-20,1)+0.06\cdot N(10,1)+0.7\cdot\delta_{0}$ and $\bm{\beta}^{\star}_{\mathsf{al}}=\sum_{i=1}^{J}\upsilon^{\star}_{i}\cdot\mathbf{o}_{\mathcal{J}(i)}$ with $\upsilon^{\star}_{i}=5\cdot\sqrt{p},i\in\{2,4,6\}$ and 00 otherwise. We set $\mathcal{J}$ to be the top 20 PCs for all designs, except MultiCauchy where we use the top 100 PCs. Penalty $h$ used in complement PCR step (described in Section 4.3.2) is identical to that used in Figure 1. See the corresponding QQ plot in Figure 6 from Appendix. bottom-left: Same setting as top-left except that the designs are taken from real datasets (real data descriptions are in Definition E.5 from Appendix). The dataset sizes are as follows. $\mathsf{Speech}$: $200\times 400$, $\mathsf{DNA}$: $100\times 180$, $\mathsf{SP500}$: $300\times 496$ and $\mathsf{FaceImage}$: $1348\times 2914$ and $\mathsf{Crime}$: $50\times 99$. All designs are centered and standardized (across rows) and then rescaled so that the average of eigenvalues of sample covariance matrix is $1$. The signal is generated in the same way as in the top-left. See the corresponding QQ plot in Figure 7 from Appendix. We set $\mathcal{J}$ to be the top 10 PCs for all designs, except FaceImage where we once again use the top 100 PCs. top-right: Under the setting of top-left, we plot the false coverage proportion (FCP) of the confidence intervals for $({\beta}^{\star}_{i})_{i=1}^{p}$, as we vary the targeted FCP level on the x-axis $\alpha$ from 00 to $1$. The $y$-axis also ranges from 0 to 1. bottom-right: analogous FCP plots under the setting of bottom-left.
@@ 4.3.2 Complement PCR
We leverage our Spectrum-Aware Debiasing theory to devise a modified PCR procedure that provides an accurate estimate of the complement component $\bm{\zeta}^{\star}$. We collect the indices in $\mathcal{J}^{c}$ corresponding to positive eigenvalues not used by Alignment PCR
@@
$\mathbf{X}_{\mathsf{new}}:=\quantity(\mathbf{D}_{\bar{\mathcal{J}}}^{\top}\mathbf{D}_{\bar{\mathcal{J}}})^{1/2}\mathbf{O}_{\bar{\mathcal{J}}},\;\;\;\;\mathbf{y}_{\mathsf{new}}:=\quantity(\mathbf{D}_{\bar{\mathcal{J}}}^{\top}\mathbf{D}_{\bar{\mathcal{J}}})^{1/2}\hat{\bm{\theta}}_{\mathsf{pcr}}^{\bot},$ (32)
where $\mathbf{D}_{\bar{\mathcal{J}}}\in\mathbb{R}^{n\times(N-J)}$, $\mathbf{O}_{\bar{\mathcal{J}}}\in\mathbb{R}^{(N-J)\times p}$ respectively consist of the columns of $\mathbf{D}$ and the rows of $\mathbf{O}$ indexed by $\bar{\mathcal{J}}$. We employ Spectrum-Aware Debiasing on this new dataset. The resulting estimator, which we call complement PCR, is $\hat{\bm{\beta}}_{\mathsf{co}}=\hat{\bm{\beta}}^{u}(\mathbf{X}_{\mathsf{new}},\mathbf{y}_{\mathsf{new}},h)$, calculated using this new dataset in (3). We establish in Theorem 4.1 (b) that $\hat{\bm{\beta}}_{\mathsf{co}}$ is approximately Gaussian centered at the complement signal component $\bm{\zeta}^{\star}$, with variance $\hat{\tau}_{*}=\hat{\tau}_{*}(\mathbf{X}_{\mathsf{new}},\mathbf{y}_{\mathsf{new}},h)$ obtained using (26) on the new dataset.

### v6

@@ 3.1 Well-definedness of our debiasing formula
Here $h^{\prime\prime}(\cdot)$ refers to the extended version we defined using Lemma 2.8 where one should plug in $h^{\prime\prime}(x)=+\infty$ if $h$ is not twice continuously differentiable at $x$.
The following Proposition is restated from Proposition C.12 in Appendix.
@@
admits a unique solution in $(0,+\infty)$ if and only if there exists some $i\in[p]$ such that $h^{\prime\prime}(\hat{{\beta}}_{i})\neq+\infty$ and at least one of the following holds: (i) $\left\|h^{\prime\prime}(\hat{\bm{\beta}})\right\|_{0}=p$; (ii) $\mathbf{X}^{\top}\mathbf{X}$ is non-singular; (iii) $\norm{d}_{0}+\norm{h^{\prime\prime}(\hat{\bm{\beta}})}_{0}>p$.
Remark 3.2.
The assumptions of Proposition 3.1 hold under Assumption 15 for all $p$ sufficiently large. See the proof of Proposition C.15. Furthermore, if $h$ is the Lasso penalty, Assumption 5 maybe dropped and the assumptions of Proposition 3.1 hold under the assumptions of Proposition H.5. See the proof of Proposition H.5.
Remark 3.3.
We emphasize that the appearance of $h^{\prime\prime}$ in (24) does not preclude interesting cases such as the Lasso or Elastic Net where the penalty is non-differentiable only on a finite set. As in Lemma 2.8, one may simply replace $h^{\prime\prime}(x_{0})$ to be $\infty$ and hence the corresponding summand term $\frac{1}{\gamma+h^{\prime\prime}(x_{0})}$ with 0 if $h$ is not differentiable at $x_{0}$. That said, using Lemma 2.8, we could also express (4) in terms of $\operatorname{Prox}_{vh}^{\prime}(x)$ for a suitable constant $v$ (instead of $h^{\prime\prime}$). The latter formulation is more common in the previous debiasing literature [12]. But in the way we have set things up, these formulations are equivalent.
@@ 3.2 The procedure
[…]
@@ 3.3 Asymptotic normality
Theorem 3.1 below states that the empirical distribution of $(\hat{\tau}_{*}^{-1/2}(\hat{{\beta}}^{u}_{i}-{\beta}^{\star}_{i}))_{i=1}^{p}$ converges to a standard Gaussian. The proof and discussion of technical novelty is deferred to Appendix C.
@@
Remark 3.6.
We prove that the asymptotic normality result in Theorem 3.1 continue to hold under a broader spectral universality class defined in Definition F.1. This result is stated in Theorem F.1.
Next, we describe a consistent estimator for the asymptotic variance $\tau_{*}$. We require some intermediate quantities that depend on the observed data and the choice of the penalty. We define these next. Later in Appendix C, we will provide intuition as to why these intermediate quantities are important and how we construct the variance estimator.
@@
$\hat{\eta}_{*}(\mathbf{X},\mathbf{y},h)\leftarrow\left(\frac{1}{p}{{\sum}}_{j=1}^{p}\frac{1}{\widehat{\mathsf{adj}}+h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)}\right)^{-1},$ (28)
$\hat{\tau}_{**}(\mathbf{X},\mathbf{y},h)\leftarrow\frac{\left\|\quantity(\mathbf{I}_{n}+\frac{1}{\hat{\eta}_{*}-\widehat{\mathsf{adj}}}\mathbf{X}\mathbf{X}^{\top})\quantity(\mathbf{y}-\mathbf{X}\hat{\bm{\beta}})\right\|^{2}-n\hat{\sigma}^{2}}{\sum_{i=1}^{p}d_{i}^{2}},$
$\hat{\tau}_{*}(\mathbf{X},\mathbf{y},h)\leftarrow\frac{1}{p}{{\sum}}_{i=1}^{p}\frac{\hat{\eta}_{*}^{2}d_{i}^{2}\hat{\sigma}^{2}+\quantity(d_{i}^{2}-\widehat{\mathsf{adj}}+2\hat{\eta}_{*})\quantity(\widehat{\mathsf{adj}}-d_{i}^{2})\quantity(\hat{\eta}_{*}-\widehat{\mathsf{adj}})^{2}\hat{\tau}_{**}}{\quantity(d_{i}^{2}-\widehat{\mathsf{adj}}+\hat{\eta}_{*})^{2}\quantity(\widehat{\mathsf{adj}})^{2}},$
@@
We illustrate Theorem 3.1 in Figure 1 under five different right-rotationally-invariant designs (cf. Remark J.2 in Appendix) with non-trivial correlation structures, and compare with Degrees-of-Freedom Debiasing with $\mathbf{M}=\mathbf{I}_{p}$. The corresponding QQ-plot can be found in Figure 5 in Appendix. We observe that our method outperforms Degrees-of-Freedom Debiasing by a margin.
@@
We call a sequence of random variables $\left(\mathsf{V}_{i}\right)_{i=1}^{p}$ exchangeable if for any permutation $\pi$ of the indices 1,,p1,...,p, the joint distribution of the permuted sequence $\left(\mathsf{V}_{\pi(i)}\right)_{i=1}^{p}$ is the same as the original sequence.
Corollary 3.10 below is a consequence of Theorem 3.1. We defer its proof to Section C.7 in Appendix. A numerical demonstration is included in Section J.3 from Appendix.
@@ 3.4 Inference
Figure 2: The above plots the TPR and FPR of the hypothesis testing procedure defined in (30) with significance level $\alpha$ and FCP of the constructed confidence intervals (31) with $b=\Phi^{-1}(1-\alpha/2),a=\Phi^{-1}(\alpha/2)$ as $\alpha$ on the x-axis varies from 0 to $1$, for both degrees-of-freedom ($\mathsf{DF}$, blue) adjustment and Spectrum-Aware ($\mathsf{SA}$, red) adjustment. The setting here is the same as in Figure 1.
@@
for any $p\geq 1$. Theorem 3.1 directly yield guarantees on the FPR, TPR and FCP as shown in Corollary 3.11 below. We defer the proof to Section C.8 in Appendix.
@@
[…]
    Then for any fixed $i$ such that ${\beta}^{\star}_{i}=0$, we have $\lim_{p\rightarrow\infty}\mathbb{P}\left(T_{i}=1\right)=\alpha,$ and the false positive rate satisfies that almost surely $\lim_{p\to\infty}\mathsf{FPR}(p)=\alpha.$ Refer also to Remark C.17 from Appendix for the exact asymptotic limit of TPR.
  • (b)
@@ 4.1 Outliers and PC alignment
The assumptions made in our previous section exclude important scenarios where the design may contain outlier eigenvalues or the signal may align with an eigenvector of the sample covariance matrix. To address this challenge, we propose an enhanced Spectrum-Aware procedure which relaxes Assumptions 1 and 2 to Assumption 8 below. To this end, denote $\mathcal{N}:=\left\{i\in[p]:d_{i}^{2}>0\right\},N:=|\mathcal{N}|$. We let $\mathcal{J}$ be a user-chosen, finite index set $\mathcal{J}\subseteq\mathcal{N}$ that should ideally contain outlier eigenvalues and alignment eigenvectors (See Remark 4.2). We denote its size as $|J:=\absolutevalue{\mathcal{J}}$.
@@
where we used $\mathcal{J}(i)$ to denote the $i$-th index in $\mathcal{J}$. Both $\bm{\upsilon}^{\star}$ and $\bm{\zeta}^{\star}$ are unknown, and they can be either deterministic or random independent of $\mathbf{O},\bm{\varepsilon}$. If $\bm{\zeta}^{\star}$ is deterministic, we assume that $\bm{\zeta}^{\star}\stackrel{{\scriptstyle W_{2}}}{{\to}}\mathsf{C}^{\star}$ as $n,p\to\infty$, where $\mathsf{C}^{\star}$ is a random variable with finite variance. If $\bm{\zeta}^{\star}$ is random, we assume the same convergence holds almost surely. Furthermore, we assume that Assumption 1 holds except that, instead of (12) and (13), we only require eigenvalues outside of $\mathcal{J}$ to be bounded and converge in empirical measure,
$\mathbf{d}_{\mathcal{J}^{c}}\stackrel{{\scriptstyle W_{2}}}{{\to}}\mathsf{D},\qquad\limsup_{p\to\infty}\max_{i\in[p]\setminus\mathcal{J}}d_{i}^{2}<+\infty$
where $\mathbf{d}_{\mathcal{J}^{c}}$ denotes a sub-vector of $\mathbf{d}=\mathbf{D}^{\top}\bm{1}_{n\times 1}$ with entries indexed by $\mathcal{J}$ removed. Finally, we require that $\limsup_{p\to\infty}\max_{i\in\mathcal{J}}d_{i}^{-2}/p\to 0.$
Under Assumption 8, $\bm{\beta}^{\star}_{\mathsf{al}}$ is the alignment component that aligns to $\mathcal{J}(i)$-th Principal Component (PC) $\mathbf{o}_{\mathcal{J}(i)}$ if the corresponding $\upsilon^{\star}_{i}$ is non-zero, while $\bm{\zeta}^{\star}$ is the complement component that is independent of the design. Note that when $\mathcal{J}=\emptyset$, Assumption 8 reduces to Assumptions 1 and 2 precisely. Finally, we note that the condition $\limsup_{p\to\infty}\max_{i\in\mathcal{J}}d_{i}^{-2}/p\to 0$ is mild: it simply requires that the smallest eigenvalues contained in $\mathcal{J}$ does not converge to 0 at a faster than $O(1/p)$ rate.
@@
$\mathcal{J}$ needs to be a finite index set that contains indices of both outlier eigenvalues and alignment eigenvectors. The outlier eigenvalues can be determined by observing the spectrum of $\mathbf{X}^{\top}\mathbf{X}$. The alignment set is generally not observed. We thus proposed an alignment test in Section 4.6 for detecting the alignments. However, we remark that eigenvectors that are both dominant and align with the signal tend to distort the debiasing procedure most severely. So often just including top few PCs in $\mathcal{J}$ can significantly improve inference.
@@ 4.2 The PCR algorithms
Figure 3: top-left: Same setting as Figure 1 except for specific changes to the design distribution parameters that lead to more challenging scenarios (see details of $\mathsf{MatrixNormal}$-$\mathsf{B}$$\mathsf{MultiCauchy}$ in Definition J.4 from Appendix). Rows 1–4 correspond to: (i) DF: Degrees-of-Freedom Debiasing as described in [13]; (ii) SA: Spectrum-Aware Debiasing as described in Definition 3.5; (iii): PCRDF: PCR-Degrees-of-Freedom Debiasing, that is, the procedure obtained from Section 4.3.3 on substituting Degrees-of-Freedom Debiasing for Spectrum-Aware Debiasing in the complement PCR step; (iv) PCRSA: PCR-Spectrum-Aware Debiasing as described in Section 4.3.3. The true signals $\bm{\beta}^{\star}$ for these designs are generated such that they align with the 2nd, 4th, and 6th PCs. Concretely, we generate the signal as follows: $\bm{\beta}^{\star}=\bm{\beta}^{\star}_{\mathsf{al}}+\bm{\zeta}^{\star}$ where the components of ${\zeta}^{\star}_{i}$ are i.i.d. draws satisfying ${\zeta}^{\star}_{i}\sim 0.24\cdot N(-20,1)+0.06\cdot N(10,1)+0.7\cdot\delta_{0}$ and $\bm{\beta}^{\star}_{\mathsf{al}}=\sum_{i=1}^{J}\upsilon^{\star}_{i}\cdot\mathbf{o}_{\mathcal{J}(i)}$ with $\upsilon^{\star}_{i}=5\cdot\sqrt{p},i\in\{2,4,6\}$ and 0 otherwise. We set $\mathcal{J}$ to be the top 20 PCs for all designs, except MultiCauchy where we use the top 100 PCs. Penalty $h$ used in complement PCR step (described in Section 4.3.2) is identical to that used in Figure 1. See the corresponding $Q$ plot in Figure 6 from Appendix. bottom-left: Same setting as top-left except that the designs are taken from real datasets (real data descriptions are in Definition J.5 from Appendix). The dataset sizes are as follows. $\mathsf{Speech}$: $200\times 400$, $\mathsf{DNA}$: $100\times 180$, $\mathsf{SP500}$: $300\times 496$ and $\mathsf{FaceImage}$: $1348\times 2914$ and $\mathsf{Crime}$: $50\times 99$. All designs are centered and standardized (across rows) and then rescaled so that the average of eigenvalues of sample covariance matrix is 11. The signal is generated in the same way as in the top-left. See the corresponding $Q$ plot in Figure 7 from Appendix. We set $\mathcal{J}$ to be the top 10 PCs for all designs, except FaceImage where we once again use the top 100 PCs. top-right: Under the setting of top-left, we plot the false coverage proportion (FCP) of the confidence intervals for $({\beta}^{\star}_{i})_{i=1}^{p}$, as we vary the targeted FCP level on the x-axis $\alpha$ from 0 to $1$. The $y$-axis also ranges from 0 to 1. bottom-right: analogous FCP plots under the setting of bottom-left.
@@ 4.3.2 Complement PCR
We leverage our Spectrum-Aware Debiasing theory to devise a modified PCR procedure that provides an accurate estimate of the complement component $\bm{\zeta}^{\star}$. We collect the indices in $\mathcal{J}^{c}$ corresponding to positive eigenvalues not used by alignment PCR ($\bar{\mathcal{J}}$ differs from $\mathcal{J}^{c}$ as it excludes zero eigenvalues)
@@
$\mathbf{X}_{\mathsf{new}}=\quantity(\mathbf{D}_{\bar{\mathcal{J}}}^{\top}\mathbf{D}_{\bar{\mathcal{J}}})^{1/2}\mathbf{O}_{\bar{\mathcal{J}}},\;\;\;\;\mathbf{y}_{\mathsf{new}}=\quantity(\mathbf{D}_{\bar{\mathcal{J}}}^{\top}\mathbf{D}_{\bar{\mathcal{J}}})^{1/2}\hat{\bm{\theta}}_{\mathsf{pcr}}^{\bot},$ (34)
where $\mathbf{D}_{\bar{\mathcal{J}}}\in\mathbb{R}^{n\times(N-J)}$, $\mathbf{O}_{\bar{\mathcal{J}}}\in\mathbb{R}^{(N-J)\times p}$ respectively consist of the columns of $\mathbf{D}$ and the rows of $\mathbf{O}$ indexed by $\bar{\mathcal{J}}$. We employ Spectrum-Aware Debiasing on this new dataset. The resulting estimator, which we call complement PCR, is $\hat{\bm{\beta}}_{\mathsf{co}}=\hat{\bm{\beta}}^{u}(\mathbf{X}_{\mathsf{new}},\mathbf{y}_{\mathsf{new}},h)$, which is calculated from (3) and (4) with respect to the new dataset $(\mathbf{X}_{\mathsf{new}},\mathbf{y}_{\mathsf{new}})$. We establish in Theorem 4.1 (b) that $\hat{\bm{\beta}}_{\mathsf{co}}$ is approximately Gaussian centered at the complement signal component $\bm{\zeta}^{\star}$, with variance $\hat{\tau}_{*}=\hat{\tau}_{*}(\mathbf{X}_{\mathsf{new}},\mathbf{y}_{\mathsf{new}},h)$ obtained using (28) on the new dataset.